OpenStudy (anonymous):

The gross income of a business over time can be represented as a set of ordered pairs in which the first element,x,is the number of the year and the second element,y, is the gross income. The gross income is given for years 1 and 2. A constant rate of change is expected for the next several years. Year 1 (1999) gross income (in 1000's)=$25 Year 2 (1998) gross income (in 1000's)=$2o The ordered pair that represents the year and gross income for 1997 is ( , ) figure out those numbers.

OpenStudy (anonymous):

the rate of change is 5 (as 25-20=5) So if you go back 1 year than it will be 20-5=15 Year 3 (1997) =25dollar

OpenStudy (anonymous):

15 I mean

OpenStudy (anonymous):

am never good at handling money

OpenStudy (anonymous):

:) well it is important to learn that

OpenStudy (anonymous):

It is two numbers in the set. Is it 25 and 151?

OpenStudy (amistre64):

you know, i got a line finder ih here someplace that would make this easier

OpenStudy (amistre64):

income = (5/1) (x -1999) +25

OpenStudy (amistre64):

the ordered pair is (1997,15)

OpenStudy (anonymous):

It wants the ordered pair that represents the year and gross income for 1997 is (?,?)